OpenStudy (anonymous):

Solve the system by any method: -4x+3y=2 5x+3=-2y

OpenStudy (anonymous):

x= -13/23 y= -2/23

OpenStudy (anonymous):

Basically, when you plot these two lines (graph) you find where they intersect. This gives you the solution to this question.

OpenStudy (anonymous):

But how did you get that?

OpenStudy (anonymous):

So, you can solve the second or first equation for any variable, which is the substitution method. You can also use the elimination method and multiply the second or first equation by -1

OpenStudy (anonymous):